OpenStudy (anonymous):

Which is the correct conversion from feet to inches? 24 in = 2 ft 3.5 ft = 35 in 45 ft = 500 in 18 ft = 216 in

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

A and D A: there are 12 inches in a foot, so 2 feet would be 2x12 which is 24. D: 12x18= 216 inches :) hope this helps

OpenStudy (whpalmer4):

Here's how to work these problems reliably: \[1 \text{ foot} = 12 \text{ inches}\]divide both sides by 12 inches: \[\frac{1 \text{ foot}}{12 \text{ inches}} = 1\] now you can multiply anything by 1 without changing its value, so you can multiply or divide with that fraction to do a conversion that will give you the correct answer. \[24 \text{ inches} * \frac{1 \text{ foot}}{12 \text{ inches}} =24\cancel{ \text{ inches} }* \frac{1 \text{ foot}}{12\cancel{ \text{ inches}}} = \frac{24 \text{ foot}}{12} = 2 \text{ foot}\]well, feet, okay :-) going the other way, \[2\text{ feet} * \frac{12 \text{ inches}}{1 \text{ foot}} = 2\cancel{\text{ feet} }* \frac{12 \text{ inches}}{1 \cancel{\text{ foot}}} = 24 \text{ inches}\] Notice that if you do it this way, you get built-in error checking. If you just haphazardly divide or multiply without using the units, you may do the wrong thing. Doing it my way, if you end up with some weird unit, you know that you've made a mistake. For example: \[2\text{ feet} * \frac{1 \text{ foot}}{12 \text{ inches}}= \frac{1 \text{ inches}^2}{12 \text{ foot}}\]and the unit makes no sense, so we've done the conversion incorrectly.